The number of levels of the independent variable is 6. How many orthogonal contrasts can be tested?

a. 1

b. 3

c. 5

d. 6
Answer #1

orthogonal contrasts can be tested = 5 (c).

[ explanation:-

number of levels of the independent variable(k) = 6

number of orthogonal contrasts = (k-1) = (6-1) = 5 ]

***in case of doubt, comment below. And if u liked the solution, please like.